OpenStudy (anonymous):

Hi, can you look at this problem? squareroot X+5 - squareroot x-3 = 4

OpenStudy (anonymous):

so x=1

OpenStudy (anonymous):

so it would be 1 +5=6

OpenStudy (anonymous):

wait hold on let me recheck

OpenStudy (anonymous):

1+5-1-3=4

OpenStudy (anonymous):

16

OpenStudy (anonymous):

ya ima pritty sure so there u go :)

OpenStudy (anonymous):

could you type your solution process?

OpenStudy (anonymous):

shira is a name of a bird in my language btw =P

OpenStudy (anonymous):

cool?

OpenStudy (anonymous):

could you explained how you solved it?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

x is seven

OpenStudy (anonymous):

7+5-7-3=16 dosent equal 16

OpenStudy (anonymous):

uggg srry im having a stupid moment i will get u better help

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@MrNood

OpenStudy (bohotness):

yes XD

OpenStudy (anonymous):

Hi, can you look at this problem? squareroot X+5 - squareroot x-3 = 4

OpenStudy (anonymous):

\[\sqrt{x+5} - \sqrt{x-3} =4\]

OpenStudy (anonymous):

yeah I dont know how to start. Should I aim to prove that threre are no solutions for x?

OpenStudy (anonymous):

thanks!!